Assume P=$500-$4Qd and our fixed cost is $1000 and the variable cost is $12 per unit.

In Excel, determine the relevant range of quantity for this problem. Create the necessary columns to determine breakeven and graph the outcome.
Answer #1

We use the excel to create a table for different quantities and find that at breakeven TR = TC.

Here TR = 500Q - 4Q^2 and TC = 1000 + 12Q. TR = TC gives Q = approximately 2 units.

The second break even occurs at Q = 119 units.

Profit is maximum at Q = 61 units when MR = MC gives 500 - 8Q = 12 or Q = 488/8 = 61 units

Quantity Price Total revenue Total cost Profit
0 500 0 1000 -1000
12 452 5424 1144 4280
24 404 9696 1288 8408
36 356 12816 1432 11384
48 308 14784 1576 13208
60 260 15600 1720 13880
72 212 15264 1864 13400
84 164 13776 2008 11768
96 116 11136 2152 8984
108 68 7344 2296 5048
120 20 2400 2440 -40

Profit maximization 18000 Profit is maximum at 61 units 16000 14000 12000 10000 Revenue and cost 8000 6000 4000 2000 o 20 40